# A User-centered Security Evaluation of Copilot

### 16 Dec 2023 v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Security and privacy Software security engineeringccs: Computing methodologies Natural language generation
@@ 1. Introduction
have the ability to become large scale producers of insecure code if left to grow unchecked. This is because are trained on code sourced from repositories that lack security guarantees, making it probable that they are trained on code segments with potential security vulnerabilities. This in turn leads to the possible production of insecure code when these are used by developers. Empirical evidence by Pearce et al. (Pearce et al., 2022) substantiates this claim, revealing that GitHub’s Copilot(GitHub Inc., 2021), a widely-used based on the Codex (Chen et al., 2021), produces insecure code approximately 40% of the time.
@@ 4.1. Participants: Recruitment and Screening
Participants for this study were recruited online via mailing lists. While our main source of participants was a university [REDACTED] computer science graduate student mailing list, we also extended invitations to industry professionals and potentially qualified undergraduate students. Participants who expressed interest in the study were asked to fill consent and screening forms which we used to determine their eligibility for the study. Selection criteria for this study was based on age (over 18 years), programming experience (at least one year of programming experience in C/C++), access to Copilot, and employment history (no affiliation with the development of Copilot, GitHub, or OpenAI). Participants who met our selection criteria were allowed to schedule a two hour online study session for the experiment to be conducted.
Like prior studies in this area, participants were not screened based on security experience because the goal of this study was to investigate the impact of Copilot use on code security among ordinary developers, regardless of security experience. We do however consider the idea of exploring the effects of participant security experience on the security of code generated with CGT assistance and interested avenue for future research.
@@ 4.2.1. Problem Design
In order to address criteria 3 and 4, we created a collection of well documented helper functions (with stub implementations) for each problem. These helper functions enabled us to expand the level of difficulty of our problems (approximating real world applications) while also constraining and guiding users towards finding solutions within a confined solution space in the allotted time. We conducted some pre-study testing to verify that Copilot could generate solutions using the stub helper functions provided in the problem file. After the study had commenced, it became apparent that there was a difference in difficulty between the two problems. We discuss this variation and how we adapted to it in section 5.1.
@@ 4.3.1. Problem Solving
We employed a within-subject study design wherein all participants served in both the treatment group and the control group. All participants solved one programming problem with Copilot access and the other problem without Copilot access. Each participant involved in the study was programmatically assigned to one of four groups on a round-robin basis. The groups determined the order in which the problems were solved and whether Copilot would be used to solve problem S or problem T. There were four groups because there were two variables that determined how participants would solve the problems, and each variable had two possibilities. Any given participant could solve problem S first or problem T first. At the same time, the participant could either solve the first problem with Copilot and the second problem without Copilot, or vice versa. Participants were given 60 minutes to solve each problem together with an instruction sheet that they could reference during problem solving. Participants were informed (verbally and in the written instructions) that they were to write secure code. There were no restrictions on the resources participants could consult to aid in solving the problem other than the restrictions on Copilot use and the use of other . Each participant’s screen was recorded during problem solving for subsequent analysis after the study session. All participants used Copilot in the Microsoft Visual Studio Code text editor. Participants were free to use and interact with Copilot in whatever manner they preferred. After each problem was finished the solutions were saved. Then, participants were required to fill out surveys to provide additional information about their perspective on the problem they just completed. Upon completion of the study, each participant was compensated CAD50.00.
@@ 4.4. Ethics
This user study obtained ethics clearance from the Human Research Ethics Board at [REDACTED]. Participant consent was obtained during the recruitment process and consenting participants were screened to ensure they met the desired criteria. Participants were informed that their screens would be recorded during the session. Data collected during sessions, including screen recordings, problem solutions, and survey information, were linked to anonymous IDs created for each participant. We maintained a key in a secure vault linking participant information (name and email address) to IDs that will be deleted once all analysis is complete and no further contact with participants is required.
@@ 5.6. Takeaways & Future work
In this study, we investigated the the impact of using Copilot on code security (RQ1) and the amount of influence that Copilot has on the presence of certain CWEs (RQ2). For the former, we observed that participants wrote more code secure for the more difficult problem when they were granted access to Copilot. In the latter, we observed no significant impact of Copilot access on the presence of any of the CWEs considered in this study.

### 03 Jan 2024 v3

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: 2024 IEEE/ACM 46th International Conference on Software Engineering; April 14–20, 2024; Lisbon, Portugalbooktitle: 2024 IEEE/ACM 46th International Conference on Software Engineering (ICSE ’24), April 14–20, 2024, Lisbon, Portugaldoi: 10.1145/3597503.3639154isbn: 979-8-4007-0217-4/24/04ccs: Security and privacy Software security engineeringccs: Computing methodologies Natural language generation
@@ 1. Introduction
have the ability to become large scale producers of insecure code if left to grow unchecked. This is because are trained on code sourced from repositories that lack security guarantees, making it probable that they are trained on code segments with security vulnerabilities. This in turn leads to the possible production of insecure code when these are used by developers. Empirical evidence by Pearce et al. (Pearce et al., 2022) substantiates this claim, revealing that GitHub’s Copilot(GitHub Inc., 2021), a widely-used based on the Codex (Chen et al., 2021), produces insecure code approximately 40% of the time.
@@ 4.1. Participants: Recruitment and Screening
Participants for this study were recruited online via mailing lists. While our main source of participants was the University of Waterloo computer science graduate student mailing list, we also extended invitations to industry professionals and potentially qualified undergraduate students. Participants who expressed interest in the study were asked to fill consent and screening forms which we used to determine their eligibility for the study. Selection criteria for this study was based on age (over 18 years), programming experience (at least one year of programming experience in C/C++), access to Copilot, and employment history (no affiliation with the development of Copilot, GitHub, or OpenAI). Participants who met our selection criteria were allowed to schedule a two hour online study session for the experiment to be conducted.
Like prior studies in this area, participants were not screened based on security experience because the goal of this study was to investigate the impact of Copilot use on code security among ordinary developers, regardless of security experience. We do however consider the idea of exploring the effects of participant security experience on the security of code generated with CGT assistance an interesting avenue for future research.
@@ 4.2.1. Problem Design
In order to address criteria 3 and 4, we created a collection of well documented helper functions (with stub implementations) for each problem. These helper functions enabled us to expand the level of difficulty of our problems (approximating real world applications) while also constraining and guiding users towards finding solutions within a confined solution space in the allotted time. We conducted pre-study testing to verify that Copilot could generate solutions using the stub helper functions provided in the problem file. After the study had commenced, it became apparent that there was a difference in difficulty between the two problems. We discuss this variation and how we adapted to it in section 5.1.
@@ 4.3.1. Problem Solving
We employed a within-subject study design wherein all participants served in both the treatment group and the control group. All participants solved one programming problem with Copilot access and the other problem without Copilot access. Each participant involved in the study was programmatically assigned to one of four groups on a round-robin basis. The groups determined the order in which the problems were solved and whether Copilot would be used to solve problem S or problem T. There were four groups because there were two variables that determined how participants would solve the problems, and each variable had two possibilities. Any given participant could solve problem S first or problem T first. At the same time, the participant could either solve the first problem with Copilot and the second problem without Copilot, or vice versa. Participants were given 60 minutes to solve each problem together with an instruction sheet that they could reference during problem solving. Participants were informed (verbally and in the written instructions) that they were to write secure code. There were no restrictions on the resources participants could consult to aid in solving the problem other than the restrictions on Copilot use and the use of other . Each participant’s screen was recorded during problem solving for subsequent analysis after the study session. All participants used Copilot in the Microsoft Visual Studio Code text editor. Participants were free to use and interact with Copilot in whatever manner they preferred. Solutions were saved once participants were done solving a given problem. Then, participants were required to fill out surveys to provide additional information about their perspective on the problem they just completed. Upon completion of the study, each participant was compensated CAD50.00.
@@ 4.4. Ethics
This user study obtained ethics clearance from the Human Research Ethics Board at the University of Waterloo. Participant consent was obtained during the recruitment process and consenting participants were screened to ensure they met the desired criteria. Participants were informed that their screens would be recorded during the session. Data collected during sessions, including screen recordings, problem solutions, and survey information, were linked to anonymous IDs created for each participant. We maintained a key in a secure vault linking participant information (name and email address) to IDs that will be deleted once all analysis is complete and no further contact with participants is required.
@@ 5.6. Takeaways & Future work
In this study, we investigated the the impact of using Copilot on code security (RQ1) and the amount of influence that Copilot has on the presence of certain CWEs (RQ2). While our findings are not statistically significant due to a limited sample size, for the former, we observed that participants wrote more code secure for the more difficult problem when they were granted access to Copilot. In the latter, we observed no significant impact of Copilot access on the presence of any of the CWEs considered in this study.